# Robust $\Lambda$-quantiles and extreme probabilities

### v1

@@ Abstract
In this paper, we investigate the robust models for $\Lambda$-quantiles with partial information regarding the loss distribution, where $\Lambda$-quantiles extend the classical quantiles by replacing the fixed probability level with a probability/loss function $\Lambda$. We find that, under some assumptions, the robust $\Lambda$-quantiles equal the $\Lambda$-quantiles of the extreme probabilities. This finding allows us to obtain the robust $\Lambda$-quantiles by applying the results of robust quantiles in the literature. Our results are applied to uncertainty sets characterized by three different constraints respectively: moment constraints, probability distance constraints via Wasserstein metric, and marginal constraints in risk aggregation. We obtain some explicit expressions for robust $\Lambda$-quantiles by deriving the extreme probabilities for each uncertainty set. Those results are applied to optimal portfolio selection under model uncertainty.
Key-words: Quantiles; $\Lambda$-quantiles; Distributionally robust optimization; Robust $\Lambda$-quantiles; Extreme probabilities; Wasserstein distance; Portfolio selection
@@ 1 Introduction
To assess the impact of model misspecification and offer a robust quantification of the optimization problem, distributionally robust optimization (DRO) has been studied extensively recently and become a fast-growing field. In DRO, some object of interest is assessed over a set of alternative distributions, where the object can be some risk measures (functionals) or expected utility and the set of alternative distributions are called the uncertainty set or ambiguity set. Typically, we are concerned with the worst-case and best-case values of some risk measures over the uncertainty set, called the robust risk measures, representing the robust quantification of the risk measures, which is independent of the choice of the distributions lying in the uncertainty set. The importance of robustness of risk measures has been emphasized in the academic response of Basel Accord 3.5 in Embrechts et al. (2014), and the robustness of the risk measures has been studied in e.g., Cont et al. (2010) and He et al. (2024).
@@
The uncertainty sets are always characterized by some constraints, representing the partially known information. The moment constraints are popularly used to characterize the uncertainty sets in the DRO literature. It means some information on the moments of the distributions is known. Under this type of uncertainty sets, robust-quantiles were studied in Ghaoui et al. (2003), robust expected shortfall (ES) was considered in Natarajan et al. (2010) and Zhu and Fukushima (2009), robust convex risk measures were investigated in Li (2018), and robust distortion risk measures and riskmetrics were considered in Pesenti et al. (2020), Shao and Zhang (2023), Shao and Zhang (2024), and Cai et al. (2023). One popular probability distance constraint is defined by Wasserstein metric, where the Wasserstein distance represents the tolerance of the discrepancy of the alternative distributions with the baseline distribution; see e.g., Esfahani and Kuhn (2018) and Blanchet and Murthy (2019). The worst-case values of quantiles, spectral risk measures and distortion riskmetrics over this uncertainty set were studied in Bernard et al. (2024), Liu et al. (2022) and Pesenti et al. (2020). The marginal constraint is always imposed on the risk aggregation, where the marginal distributions are known but the dependence structure is completely unknown. For this uncertainty set, the robust-quantiles are the main concern in the literature, which is only solved for some special marginal distributions. The robust-quantiles were derived for marginal distributions with monotone densities in Wang and Wang (2016), Jakobsons et al. (2016) and Blanchet et al. (2020). We refer to Puccetti and Rüschendorf (2012) and Eckstein et al. (2020) for numerical studies. The uncertainty sets characterized by the above three constraints will be explored later.
In this paper, we focus on the $\Lambda$-quantiles, which is an extension of quantiles by replacing the fixed probability level by a probability/loss function. As one of the most important measures of risk in quantitative risk management, quantiles have been criticized by two major reasons: First, it cannot capture the tail risk; Second, it is lack of convexity (or subadditivity), resulting in penalizing diversification. More detailed discussions on it can be found in, e.g., McNeil et al. (2015). To overcome some drawbacks of quantiles, $\Lambda$-quantiles was introduced by Frittelli et al. (2014) as an extension of quantiles. In the recent literature, it shows that $\Lambda$-quantiles possess some desirable properties. From the definition of $\Lambda$-quantiles, it is able to control the tail probability of the risk for increasing $\Lambda$ functions, which was also tested empirically in Hitaj et al. (2018). Moreover, as shown in Han et al. (2024), $\Lambda$-quantiles with decreasing $\Lambda$ functions satisfy quasi-star-shapedness, indicating that convex combination with a determinate risk does not increase the risk, which is a property weaker than quasi-convexity and penalizing the special type of concentration. In addition, $\Lambda$-quantiles with decreasing $\Lambda$ functions also satisfy cash-subadditivity, which is desirable to measure the risk of the financial position in the presence of stochastic or ambiguous interest rates; see El Karoui and Ravanelli (2009) and Han et al. (2024). We refer to Burzoni et al. (2017) for some other properties of $\Lambda$-quantiles such as robustness, elicitability, and consistency.
@@
Typically, the $\Lambda$ function is chosen to be either increasing or decreasing, representing the risk appetite of the decision maker (DM) as shown in Frittelli et al. (2014), and the choice and estimation of the $\Lambda$ function were studied in Hitaj et al. (2018). Recently, an axiomatization of $\Lambda$-quantiles was offered in Bellini and Peri (2022) and a new expression of $\Lambda$-quantiles with decreasing $\Lambda$ function was provided in Han et al. (2024). Additionally, Ince et al. (2022) focused on the risk contribution (sensitivity analysis) of $\Lambda$-quantiles, Balbás et al. (2023) studied the applications of $\Lambda$-quantiles to reinsurance contracts and premium calculation, Liu (2024) investigated the risk sharing problem among multiple agents using $\Lambda$-quantiles to represent their preferences, and Pesenti and Vanduffel (2024) used the scoring function of $\Lambda$-quantiles for elicitability as the cost function to study the Monge-Kantorovich optimal transport.
@@
In this paper, we aim to establish the robust models for $\Lambda$-quantiles with limited information on the underlying distribution of loss, where the function $\Lambda$ can be either decreasing or increasing. Despite the prevailing notion in existing literature that most researchers consider the risk in the worst-case scenario, wherein the DM is extremely ambiguity-averse, empirical studies (see e.g., Heath and Tversky (1991) and Kocher et al. (2015)) demonstrate that the DM’s attitude towards ambiguity is not uniformly negative. In fact, they may exhibit ambiguity-loving tendencies if they perceive themselves as knowledgeable or competent in the context. Thus, both the worst and best possible values for $\Lambda$-quantiles over a set of alternative distributions will be considered.
@@
In Section 3, we show that, under some assumption, the robust $\Lambda$-quantiles over a general uncertainty set is equal to the $\Lambda$-quantile of the extreme probabilities over the same uncertainty set, which is the main finding of this paper. This is an unexpected result. This fact shows that the computation of robust $\Lambda$-quantiles can be decomposed as two steps: i) compute the extreme probabilities; ii) compute the $\Lambda$-quantiles of the obtained extreme probabilities. It means that in order to compute the robust $\Lambda$-quantiles, it suffices to compute the extreme probabilities, which can be derived using the results of robust quantiles. The crucial assumption for our results is the attainability of extreme probabilities. Verifying this property for popular uncertainty sets is not trivial, and will be discussed in Section 4. Moreover, we construct several counterexamples for the assumptions of our main results and discuss the existence of the optimal distributions in the robust $\Lambda$-quantiles. It is worth mentioning that our findings are closely liked to Mao et al. (2024), where a novel approach, called model aggregation approach, was introduced to evaluate the risk under model uncertainty for some risk measures including quantiles and ES.
In Section 4, we consider the uncertainty sets characterized by three different constraints respectively: moment constraints, probability constraints via Wasserstein distance, and marginal constraints in risk aggregation. We obtain the extreme probabilities for each uncertainty set replying on the results on robust quantiles over the same uncertainty sets in the literature and check the attainability of the extreme probabilities for each uncertainty set. Applying our main results in Section 3, we directly obtain the robust $\Lambda$-quantiles, showing the usefulness of our findings. Moreover, we apply our results to the portfolio selection problem with model uncertainty, where the risk of the portfolio is assessed by $\Lambda$-quantiles. In particular, for the uncertainty sets with marginal constraints in risk aggregation, we show that diversification is penalized for the portfolio of assets with the same marginal distribution, extending the result for quantiles in Chen et al. (2022). Some numerical studies have been done to show the values of the robust $\Lambda$-quantiles and optimal portfolio positions. The paper is concluded in Section 5.
@@ 2 Notation and preliminaries
Let $\mathcal{X}$ be a set of random variables in a given atomless probability space $(\Omega,\mathcal{F},\mathbb{P})$. Denote by $\mathcal{M}$ the set of cumulative distribution functions (cdfs) of all random variables in $\mathcal{X}$. For any law-invariant risk measures, it can be defined on either $\mathcal{X}$ or $\mathcal{M}$, which makes no difference for the understanding. For a distribution $F\in\mathcal{M}$, its left quantile at level $\alpha\in[0,1]$ is defined by
@@
with the convention that $\inf\emptyset=\infty$. Note that by the above definition, we have $q_{0}^{-}(F)=-\infty$ and $q_{1}^{+}(F)=\infty$. We refer to e.g., McNeil et al. (2015) for more discussions on the properties and application of quantiles. We next present four definitions of $\Lambda$-quantiles given in Bellini and Peri (2022).
@@
In Definition 1, agents can choose $\Lambda$ as either increasing or decreasing functions. Throughout this paper, “increasing” (resp. “decreasing”) means non-decreasing (resp. non-increasing). Without additional assumptions, the four $\Lambda$-quantiles can all differ; see Examples 4 and 5 of Bellini and Peri (2022). In general, an increasing $\Lambda$ implies that the DM only accepts smaller probability for larger losses. This reflects a relatively more risk-averse attitude. Conversely, a decreasing $\Lambda$ suggests a tendency to tolerate a larger probability as losses increase. This indicates that the DM is more relatively risk-seeking.
A simple example of $\Lambda$-quantile is the two-level $\Lambda$-quantile presented in Example 7 of Bellini and Peri (2022), where the $\Lambda$ function is defined as $\Lambda:x\mapsto\beta\mathds{1}_{\{x<z\}}+\alpha\mathds{1}_{\{x\geqslant z\}}$. In particular, if $\Lambda$ is a constant, then $\Lambda$-quantiles boil down to the quantiles.
A mapping $\rho:\mathcal{X}\rightarrow\mathbb{R}$ is said to satisfy monotonicity if $\rho(X)\leqslant\rho(Y)$ for all $X,Y\in\mathcal{X}$ with $X\leqslant Y$, and satisfy cash additivity if $\rho(X+m)=\rho(X)+m$ for all $X\in\mathcal{X}$ and $m\in\mathbb{R}$. It is well known that $q^{-}_{\alpha}$ and $q^{+}_{\alpha}$ satisfy the both two properties. However, $\Lambda$-quantiles are not cash-additive in general.
A mapping $\rho:\mathcal{X}\rightarrow\mathbb{R}$ is said to satisfy cash subadditivity (resp. cash supadditivity) if $\rho(X+m)\leqslant\rho(X)+m$ (resp. $\rho(X+m)\geqslant\rho(X)+m$) for all $X\in\mathcal{X}$ and $m\geqslant 0$. A mapping $\rho$ is said to be quasi-star-shapedness if $\rho(\lambda X+(1-\lambda)t)\leqslant\max\{\rho(X),\rho(t)\}$ for all $X\in\mathcal{X},t\in\mathbb{R}$ and $\lambda\in[0,1]$. In the context of capital requirement, cash subadditivity or supadditivity allows a non-linear increase of the capital requirement as cash is added to the future financial position. Cash-subadditivity is desirable if the interest rate is stochastic or ambiguous, and it allows the current reserve and the future risk positions to preserve their own numéraire. Here cash-subadditivity reflects the time value of money; see El Karoui and Ravanelli (2009) and Han et al. (2024). Quasi-star-shapedness means that $\rho$ has quasi-convexity at each constant, representing some consideration of diversification, which is weaker than quasi-convexity. Its theoretic-decision interpretation is available at Han et al. (2024).
@@
Note that (2) follows directly from Theorem 3.1 of Han et al. (2024). Next, we show (3). Note that for $x\in\mathbb{R}$ and $t\in[0,1]$, $F(x)\leqslant t$ implies $q^{+}_{t}(F)\geqslant x$. Hence, for a decreasing $\Lambda$, we have
@@
Moreover, using the fact $F(x)>t$ implies $q^{+}_{t}(F)\leqslant x$, we have
@@
In Bellini and Peri (2022), many properties of $\Lambda$-quantiles are discussed such as monotonicity, locality, and quasi-convexity on distributions. We next summarize some of the properties possessed by $\Lambda$-quantiles that were not included in Bellini and Peri (2022) in the following proposition.
@@
The following statements hold.
@@
    $q_{\Lambda}^{-}(F)\leqslant\tilde{q}_{\Lambda}^{-}(F)$ and $q_{\Lambda}^{+}(F)\leqslant\tilde{q}_{\Lambda}^{+}(F)$, and the inequalities become equalities if $\Lambda$ is decreasing;
@@
    All the four $\Lambda$-quantiles $q_{\Lambda}^{-},q_{\Lambda}^{+},\tilde{q}_{\Lambda}^{-}$, and $\tilde{q}_{\Lambda}^{+}$ are cash-subadditive if $\Lambda$ is decreasing and cash-supadditive if $\Lambda$ is increasing;
@@
    If $\Lambda$ is decreasing, all the four $\Lambda$-quantiles $q_{\Lambda}^{-},q_{\Lambda}^{+},\tilde{q}_{\Lambda}^{-}$, and $\tilde{q}_{\Lambda}^{+}$ are quasi-star-shaped.
@@
We first focus on (i). We first show $q_{\Lambda}^{-}(F)\leqslant\tilde{q}_{\Lambda}^{-}(F)$. Clearly, if $\tilde{q}_{\Lambda}^{-}(F)=-\infty$, then $F(x)\geqslant\Lambda(x)$ for all $x\in\mathbb{R}$. This implies $q_{\Lambda}^{-}(F)=-\infty$. Hence, $q_{\Lambda}^{-}(F)\leqslant\tilde{q}_{\Lambda}^{-}(F)$ holds. If $\tilde{q}_{\Lambda}^{-}(F)=+\infty$, the inequality holds trivially. Next, we consider the case $x_{0}=\tilde{q}_{\Lambda}^{-}(F)\in\mathbb{R}$. By the definition, we have $F(x)\geqslant\Lambda(x)$ for all $x>x_{0}$, which implies $q_{\Lambda}^{-}(F)\leqslant x_{0}=\tilde{q}_{\Lambda}^{-}(F)$. We can similarly show $q_{\Lambda}^{+}(F)\leqslant\tilde{q}_{\Lambda}^{+}(F)$. The equalities for decreasing $\Lambda$ follows directly from Proposition 6 of Bellini and Peri (2022).

### v2

# Robust $\Lambda$-quantiles and extremal distributions

@@ Abstract
In this paper, we investigate the robust models for $\Lambda$-quantiles with partial information regarding the loss distribution, where $\Lambda$-quantiles extend the classical quantiles by replacing the fixed probability level with a probability/loss function $\Lambda$. We find that, under some assumptions, the robust $\Lambda$-quantiles equal the $\Lambda$-quantiles of the extremal distributions. This finding allows us to obtain the robust $\Lambda$-quantiles by applying the results of robust quantiles in the literature. Our results are applied to uncertainty sets characterized by three different constraints respectively: moment constraints, probability distance constraints via the Wasserstein metric, and marginal constraints in risk aggregation. We obtain some explicit expressions for robust $\Lambda$-quantiles by deriving the extremal distributions for each uncertainty set. These results are applied to optimal portfolio selection under model uncertainty.
Key-words: Quantiles; $\Lambda$-quantiles; Distributionally robust optimization; Robust $\Lambda$-quantiles; Extremal distributions; Wasserstein distance; Portfolio selection.
@@ 1 Introduction
To assess the impact of model misspecification and offer a robust quantification of the optimization problem, distributionally robust optimization (DRO) has been extensively studied in recent years and has become a fast-growing field. In DRO, some object of interest is assessed over a set of alternative distributions, where the object can be some risk measures (functionals) or expected utility and the set of alternative distributions is called the uncertainty set or ambiguity set. Typically, we are concerned with the “worst-case” (maximal) and “best-case” (minimal) values of some risk measures over the uncertainty set, called the robust risk measures, representing the robust quantification of the risk measures, which is independent of the choice of the distributions lying in the uncertainty set. The importance of the robustness of risk measures has been emphasized in the academic response of Basel Accord 3.5 in Embrechts et al. (2014), and the robustness of the risk measures has been studied in e.g., Cont et al. (2010) and He et al. (2024).
@@
The uncertainty sets are always characterized by some constraints, representing the partially known information. The moment constraints are popularly used to characterize the uncertainty sets in the DRO literature. This means that some information on the moments of the distributions is known. Under this type of uncertainty sets, robust-quantiles were studied in Ghaoui et al. (2003), robust expected shortfall (ES) was considered in Natarajan et al. (2010) and Zhu and Fukushima (2009), robust convex risk measures were investigated in Li (2018), and robust distortion risk measures and riskmetrics were considered in Pesenti et al. (2024), Shao and Zhang (2023), Shao and Zhang (2024), and Cai et al. (2023). One popular probability distance constraint is defined by the Wasserstein metric, where the Wasserstein distance represents the tolerance of the discrepancy of the alternative distributions with the baseline distribution; see e.g., Esfahani and Kuhn (2018) and Blanchet and Murthy (2019). The worst-case values of quantiles, spectral risk measures and distortion riskmetrics over this uncertainty set were studied in Liu et al. (2022), Bernard et al. (2024) and Pesenti et al. (2024). The marginal constraint is always imposed on the risk aggregation, where the marginal distributions are known but the dependence structure is completely unknown. For this uncertainty set, the robust-quantiles are the main concern in the literature, which is only solved for some special marginal distributions. The robust-quantiles were derived for marginal distributions with monotone densities in Wang and Wang (2016), Jakobsons et al. (2016) and Blanchet et al. (2024). We refer to Puccetti and Rüschendorf (2012) and Eckstein et al. (2020) for numerical studies. The uncertainty sets characterized by the above three constraints will be explored later. Recently, the robust models for quantiles under ambiguity were obtained in Liu et al. (2024), where the uncertainty set is a collection of probability measures, different from our setup in the paper.
In this paper, we focus on $\Lambda$-quantiles, first introduced by Frittelli et al. (2014), which extend the classical quantiles by replacing the fixed probability level with a probability/loss function. For $\Lambda:\mathbb{R}\to[0,1]$ and a distribution function $F$, one version of $\Lambda$-quantiles is defined as
$q_{\Lambda}^{-}(F)=\inf\{x\in\mathbb{R}:F(x)\geqslant\Lambda(x)\}$
with the convention that $\inf\emptyset=\infty$. Four different definitions of $\Lambda$-quantiles introduced in Bellini and Peri (2022) will be given in Section 2.
The $\Lambda$-quantiles offer greater flexibility by choosing the function $\Lambda$ instead of fixing a confidence level, which allows one to handle multiple confidence levels simultaneously. Typically, the $\Lambda$ function is chosen to be either increasing or decreasing, and the choice and estimation of the $\Lambda$ function were studied in Hitaj et al. (2018). Interestingly, $\Lambda$-quantiles still retain some good properties of quantiles such as monotonicity, robustness, elicitability, consistency and locality; see Burzoni et al. (2017) and Bellini and Peri (2022). As shown in Han et al. (2025), $\Lambda$-quantiles with decreasing $\Lambda$ functions satisfy quasi-star-shapedness, indicating that convex combination with a deterministic risk does not increase the risk, which is a property weaker than quasi-convexity and penalizing the special type of concentration. In addition, $\Lambda$-quantiles with decreasing $\Lambda$ functions also satisfy cash-subadditivity, which is desirable for measuring the risk of the financial position in the presence of stochastic or ambiguous interest rates; see El Karoui and Ravanelli (2009) and Han et al. (2025). Recently, an axiomatization of $\Lambda$-quantiles was offered in Bellini and Peri (2022), and a new expression of $\Lambda$-quantiles with a decreasing $\Lambda$ function was provided in Han et al. (2025).
@@
The application of $\Lambda$-quantiles has also attracted some attention recently. For instance, Ince et al. (2022) studied the risk contribution (sensitivity analysis) of $\Lambda$-quantiles; Balbás et al. (2023) and Boonen et al. (2025) focused on the applications of $\Lambda$-quantiles to optimal reinsurance design and premium calculation; Liu (2024) and Xia and Hu (2024) investigated the risk-sharing problem with $\Lambda$-quantiles representing the agents’ preferences, which was extended to the case with heterogeneous beliefs by Liu et al. (2024); and Pesenti and Vanduffel (2024) used the scoring function of $\Lambda$-quantiles for elicitability as the cost function to study the Monge-Kantorovich optimal transport.
@@
In this paper, we aim to establish the robust models for $\Lambda$-quantiles with limited information on the underlying distribution of loss, where the function $\Lambda$ can either decrease or increase. Despite the prevailing notion in existing literature that most researchers consider the risk in the worst-case scenario, wherein the decision maker (DM) is extremely ambiguity-averse, empirical studies (see e.g., Heath and Tversky (1991) and Kocher et al. (2015)) demonstrate that the DM’s attitude toward ambiguity is not uniformly negative. In fact, they may exhibit ambiguity-loving tendencies if they perceive themselves as knowledgeable or competent in the context. Thus, both the worst and best possible values for $\Lambda$-quantiles over a set of alternative distributions will be considered.
@@
In Section 3, we show that, under some assumption, the robust $\Lambda$-quantiles over a general uncertainty set is equal to the $\Lambda$-quantile of the extremal distributions over the same uncertainty set, which is the main finding of this paper. In this paper, the extremal distribution is an increasing function representing the pointwise supremum or infimum of the distribution functions over an uncertainty set. This fact shows that the computation of robust $\Lambda$-quantiles can be decomposed into two steps under some conditions: i) compute the extremal distributions, and ii) compute the $\Lambda$-quantiles of the obtained extremal distributions. This means that, under some conditions, in order to compute the robust $\Lambda$-quantiles, it suffices to compute the extremal distributions, which can be derived via the results of robust quantiles. The crucial assumption for our results is the attainability of extremal distributions. Verifying this property for popular uncertainty sets is not trivial, and will be discussed in Section 4. Moreover, we construct several counterexamples for the assumptions of our main results and discuss the existence of the optimal distributions in the robust $\Lambda$-quantiles. Notably, our findings are closely linked to Mao et al. (2024), where a novel approach, called the model aggregation approach, was introduced to evaluate the risk under model uncertainty for some risk measures including quantiles and ES.
In Section 4, we consider the uncertainty sets characterized by three different constraints respectively: moment constraints, probability constraints via the Wasserstein metric, and marginal constraints in risk aggregation. We obtain the extremal distributions for each uncertainty set replying on the results on robust quantiles over the same uncertainty sets in the literature and check the attainability of the extremal distributions for each uncertainty set. By applying our main results in Section 3, we directly obtain the robust $\Lambda$-quantiles, showing the usefulness of our findings. Moreover, we apply our results to the portfolio selection problem with model uncertainty, where the risk of the portfolio is assessed by $\Lambda$-quantiles. In particular, for the uncertainty sets with marginal constraints in risk aggregation, we show that diversification is penalized for the portfolio of assets with the same marginal distribution, extending the results for quantiles in Chen et al. (2022). Some numerical studies have been performed to show the values of the robust $\Lambda$-quantiles and optimal portfolio positions. The paper is concluded in Section 5.
@@ 2 Notation and preliminaries
Let $\mathcal{X}$ be a set of random variables on a given atomless probability space $(\Omega,\mathcal{F},\mathbb{P})$. Denote by $\mathcal{M}$ the set of cumulative distribution functions (cdfs) of all random variables in $\mathcal{X}$. In this paper, we consider the risk functional $\rho:\mathcal{M}\to\mathbb{R}$. For convenience, we sometimes use $\rho(X)$ for $X\in\mathcal{X}$ to represent $\rho(F_{X})$, where $F_{X}$ is the distribution of $X$. For a distribution $F\in\mathcal{M}$, its left quantile at level $\alpha\in[0,1]$ is defined as
@@
with the convention that $\inf\emptyset=\infty$. Note that by the above definition, we have $q_{0}^{-}(F)=-\infty$ and $q_{1}^{+}(F)=\infty$. We refer to e.g., McNeil et al. (2015) for more discussions on the properties and applications of quantiles. Throughout this paper, terms such as “increasing” and “decreasing” are in the non-strict sense, and the notation $a\vee b$ (resp. $a\wedge b$) is the maximum (resp. minimum) between real numbers $a$ and $b$, and $a_{+}=a\vee 0$.
We next present four definitions of $\Lambda$-quantiles given in Bellini and Peri (2022).
@@
In Definition 1, agents can choose $\Lambda$ as either increasing or decreasing functions. Without additional assumptions, the four $\Lambda$-quantiles can all differ; see Examples 4 and 5 of Bellini and Peri (2022). In general, an increasing $\Lambda$ implies that the DM accepts only a smaller probability for larger losses. Conversely, a decreasing $\Lambda$ suggests a tendency to tolerate a greater probability as losses increase.
A simple example of $\Lambda$-quantile is the two-level $\Lambda$-quantile presented in Example 7 of Bellini and Peri (2022), where the $\Lambda$ function is defined as $\Lambda:x\mapsto\beta\mathds{1}_{\{x<z\}}+\alpha\mathds{1}_{\{x\geqslant z\}}$. In particular, if $\Lambda$ is a constant, then $\Lambda$-quantiles boil down to quantiles.
A mapping $\rho:\mathcal{X}\rightarrow\mathbb{R}$ is said to satisfy monotonicity if $\rho(X)\leqslant\rho(Y)$ for all $X,Y\in\mathcal{X}$ with $X\leqslant Y$, and satisfy cash-additivity if $\rho(X+m)=\rho(X)+m$ for all $X\in\mathcal{X}$ and $m\in\mathbb{R}$. It is well known that $q^{-}_{\alpha}$ and $q^{+}_{\alpha}$ satisfy both properties. However, $\Lambda$-quantiles are not cash-additive in general.
A mapping $\rho:\mathcal{X}\rightarrow\mathbb{R}$ is said to satisfy cash-subadditivity (resp. cash-supadditivity) if $\rho(X+m)\leqslant\rho(X)+m$ (resp. $\rho(X+m)\geqslant\rho(X)+m$) for all $X\in\mathcal{X}$ and $m\geqslant 0$. A mapping $\rho$ is said to be quasi-star-shaped if $\rho(\lambda X+(1-\lambda)t)\leqslant\max\{\rho(X),\rho(t)\}$ for all $X\in\mathcal{X},t\in\mathbb{R}$ and $\lambda\in[0,1]$. In the context of capital requirements, cash-subadditivity or supadditivity allows a non-linear increase in the capital requirement as cash is added to the future financial position. Cash-subadditivity is desirable if the interest rate is stochastic or ambiguous, and it allows the current reserve and the future risk positions to preserve their own numéraire. Here cash-subadditivity reflects the time value of money; see El Karoui and Ravanelli (2009) and Han et al. (2025). Cash-subadditivity is satisfied by convex loss-based risk measures proposed in Cont et al. (2013). Quasi-star-shapedness means that $\rho$ has quasi-convexity at each constant, representing some consideration of diversification, which is weaker than quasi-convexity. Its theoretic-decision interpretation is available at Han et al. (2025).
@@
Note that (2) follows directly from Theorem 3.1 of Han et al. (2025). Next, we show (3). Note that for $x\in\mathbb{R}$ and $t\in[0,1]$, $F(x)\leqslant t$ implies that $q^{+}_{t}(F)\geqslant x$. Hence, for a decreasing $\Lambda$, we have
@@
Moreover, because $F(x)>t$ implies that $q^{+}_{t}(F)\leqslant x$, we have
@@
In Bellini and Peri (2022), many properties of $\Lambda$-quantiles such as monotonicity, locality, and quasi-convexity on distributions are discussed. We next summarize some of the properties possessed by $\Lambda$-quantiles that were not included in Bellini and Peri (2022) in the following proposition.
@@
For $F\in\mathcal{M}$, the following statements hold.
@@
    $q_{\Lambda}^{-}(F)\leqslant\tilde{q}_{\Lambda}^{-}(F)$ and $q_{\Lambda}^{+}(F)\leqslant\tilde{q}_{\Lambda}^{+}(F)$, and the inequalities become equalities if $\Lambda$ is decreasing.
@@
    All four $\Lambda$-quantiles $q_{\Lambda}^{-},q_{\Lambda}^{+},\tilde{q}_{\Lambda}^{-}$, and $\tilde{q}_{\Lambda}^{+}$ are cash-subadditive if $\Lambda$ is decreasing and cash-supadditive if $\Lambda$ is increasing. Furthermore, if $\Lambda:\mathbb{R}\to(0,1)$ is right-continuous and has at most finite discontinuous points, then the converse conclusion also holds.
@@
    If $\Lambda$ is decreasing, all four $\Lambda$-quantiles $q_{\Lambda}^{-},q_{\Lambda}^{+},\tilde{q}_{\Lambda}^{-}$, and $\tilde{q}_{\Lambda}^{+}$ are quasi-star-shaped.
@@
We first focus on (i). We first show $q_{\Lambda}^{-}(F)\leqslant\tilde{q}_{\Lambda}^{-}(F)$. Clearly, if $\tilde{q}_{\Lambda}^{-}(F)=-\infty$, then $F(x)\geqslant\Lambda(x)$ for all $x\in\mathbb{R}$. This implies that $q_{\Lambda}^{-}(F)=-\infty$. Hence, $q_{\Lambda}^{-}(F)\leqslant\tilde{q}_{\Lambda}^{-}(F)$ holds. If $\tilde{q}_{\Lambda}^{-}(F)=+\infty$, the inequality holds trivially. Next, we consider the case $x_{0}=\tilde{q}_{\Lambda}^{-}(F)\in\mathbb{R}$. By the definition, we have $F(x)\geqslant\Lambda(x)$ for all $x>x_{0}$, which implies that $q_{\Lambda}^{-}(F)\leqslant x_{0}=\tilde{q}_{\Lambda}^{-}(F)$. We can similarly show $q_{\Lambda}^{+}(F)\leqslant\tilde{q}_{\Lambda}^{+}(F)$. The equalities for decreasing $\Lambda$ follow directly from Proposition 6 of Bellini and Peri (2022).